# Air-Ground Collaborative Robots for Fire and Rescue Missions: Towards Mapping and Navigation Perspective

### v1

@@ VII Conclusion and Outlook
2) Multi-modal data fusion for UAV mapping. Fire and rescue environments are often complex and changeable, and are filled with smoke, flames, obstructions, etc., which pose challenges to UAV mapping. To address such challenges, multi-sensor SLAM systems have demonstrated high accuracy and robustness in complex environments [145, 146, 147], and can effectively overcome the impact of smoke and lighting changes on visual perception [148]. To this end, the information of multiple sensors such as thermal imaging, visible light, and infrared can be integrated to perform multi-modal data fusion and UAV mapping. Such data fusion can effectively enhance the perception and mapping capabilities of UAVs, and also provide a more comprehensive description of the environment for UGVs. Therefore, the UAV mapping system with multi-modal data fusion will be an important research direction for air-ground collaborative robots in the fire and rescue environments.
@@
4) Embodied artificial intelligence (AI)-driven UGV navigation. The combination of embodied AI and robotics technology has brought profound impact on robot navigation [151]. And some work has verified that the integration with embedded AI, especially large language models (LLMs) and visual language models (VLMs), significantly enhances the robot’s scene understanding, path planning and autonomous navigation capabilities in complex environments [152, 153]. In fire and rescue missions, the combination of collaborative navigation systems and embodied AI can, on the one hand, improve the performance of UAV/UGV real-time fire scene understanding, task area identification and feature matching for UGV navigation, and on the other hand, effectively enhance the navigation performance and mission execution capabilities of UGV. This will play an integral role in the research of air-ground cooperative robot navigation.

### v2

@@ VII Conclusion and Outlook
2) Multi-modal data fusion for UAV mapping. Fire and rescue environments are often complex and changeable, and are filled with smoke, flames, obstructions, etc., which pose challenges to UAV mapping. To address such challenges, multi-sensor SLAM systems have demonstrated high accuracy and robustness in complex environments [143, 144], and can effectively overcome the impact of smoke and lighting changes on visual perception [145]. To this end, the information of multiple sensors such as thermal imaging, visible light, and infrared can be integrated to perform multi-modal data fusion and UAV mapping. Such data fusion can effectively enhance the perception and mapping capabilities of UAVs, and also provide a more comprehensive description of the environment for UGVs. Therefore, the UAV mapping system with multi-modal data fusion will be an important research direction for air-ground collaborative robots in the fire and rescue environments.
@@
4) Embodied artificial intelligence (AI)-driven UGV navigation. The combination of embodied AI and robotics technology has brought profound impact on robot navigation [148]. And some work has verified that the integration with embedded AI, especially large language models (LLMs) and visual language models (VLMs), significantly enhances the robot’s scene understanding, path planning and autonomous navigation capabilities in complex environments [149, 150, 151]. In fire and rescue missions, the combination of collaborative navigation systems and embodied AI can, on the one hand, improve the performance of UAV/UGV real-time fire scene understanding, task area identification and feature matching for UGV navigation, and on the other hand, effectively enhance the navigation performance and mission execution capabilities of UGV. This will play an integral role in the research of air-ground cooperative robot navigation.